OpenStudy (anonymous):

whats the pattern; 1,3,6,10,15

OpenStudy (anonymous):

it is increasing 2,3,4,5

OpenStudy (anonymous):

1+2=3 3+3=6 6+4=10 10+5=15

OpenStudy (anonymous):

1to 3 increases by 2 it increases next by 3 4,5,6, and so on see the pattern! increases by 1